# Actor-Critic Model Predictive Control: Differentiable Optimization meets Reinforcement Learning

### v7

@@ III-B Tracking Model Predictive Control
The tracking MPC formulation also relies on the assumption that a trajectory represents a well-posed, quadratic proxy objective for solving the end task. While this assumption holds for many practical use cases, it becomes limiting in situations where the system operates near its performance limits, such as aggressive maneuvers or high-speed tasks, where the trajectory may no longer serve as an effective proxy. In such cases, the tracking objective can result in suboptimal performance or even task failure [14].
@@ III-D General Quadratic MPC formulation
In this paper, we propose to directly search for the matrix coefficients of Eq. (4) (i.e., $Q_{k}$ and $p_{k}$) using reinforcement learning and differentiable MPC [26]. This way, we can encode the task as a general, non-differentiable reward function while at the same time maintaining computational tractability.
@@ III-E Actor-Critic Reinforcement Learning
where $\tau$ denotes a trajectory, $r(s_{k},a_{k})$ the reward function, and $\gamma\in[0,1)$ is a discount factor that discounts future rewards. Unlike real-time optimization in methods like MPC, RL can directly optimize a non-differentiable, sparse reward. Additionally, policy gradient optimization is typically performed offline through interactions with simulation. Once trained, the policy $\pi_{\theta}$ enables the computation of control signals by simply evaluating the function $a^{\ast}=\pi_{\theta}(s_{k})$ at each time step, reducing the computational complexity during deployment. For Actor-Critic reinforcement learning, the key idea is to simultaneously learn a state-value function $V_{\omega}(s)$ and learn a policy function $\pi_{\theta}$, where the value function (Critic) and policy (Actor) are parameterized by $\omega$ and $\theta$ separately. The policy is updated via the policy gradient theorem [116],
@@ III-F Actor-Critic Model Predictive Control
This paper proposes an Actor-Critic MPC controller architecture where the MPC is differentiable [26] and the cost function is learned end-to-end using RL. The MPC block is introduced as the differentiable module of the actor in an actor-critic pipeline that uses the proximal policy optimization (PPO) algorithm [117]. Our architecture is shown in Fig. 1. In contrast to previous work [118, 62], where the MPC is taken as a black-box controller and the gradient is sampled, in our case the gradient of the cost function with respect to the solution is analytically computed and propagated using a differentiable MPC [26]. Therefore, for every backward and forward pass of the actor network, we need to solve an optimization problem. Instead of resorting to task-specific engineering of the cost function, we propose a neural cost map where the $Q_{k}$ and $p_{k}$ terms are the output of the neural network layers preceding the differentiable MPC block. This allows us to encode the end task directly as a reward function, which is then trainable end-to-end using the PPO training scheme. The main benefit of this approach with respect to training a standard Multi-Layer Perceptron (MLP) end-to-end is that the final module of the actor is a model-based MPC controller,
@@
and therefore it retains its generalizability and robustness properties. The model-based controller placed after the last layer of the actor network ensures that the commands are always feasible for the dynamics at hand, and that they respect the system constraints. This differentiable MPC supports control input constraints but not state constraints, which is why we add $\bm{u}_{k}\in\mathcal{U}$ in (1). To allow for exploration, during training the control inputs are sampled from a Gaussian distribution where the mean is the output of the MPC block, and the variance is controlled by the PPO algorithm. However, during deployment the output from the MPC is used directly on the system without further sampling, retaining all properties of a model-based controller.
@@ III-G Neural Cost Map
The purpose of the diagonalization of the $Q$ matrix is to reduce the dimensionality of the learnable parameter space. Therefore, the dimensionality of the output dimension of the Cost Map is $2T(n_{state}+n_{input})$. In order to ensure the positive semi-definiteness of the $Q$ matrix and the positive definiteness of the R matrix, a lower bound on the value of these coefficients needs to be set. To this end, the last layer of the neural cost map has been chosen to be a sigmoid which allows for upper and lower bounds on the output value. These lower and upper limits are chosen equal for $Q$ and $p$, of 0.1 and 100000.0, respectively. Therefore, the final neural cost map consists of two hidden layers of width 512 with ReLUs in between and a sigmoid non-linearity at the end. The critic network consists also of two hidden layers of width 512 and ReLUs. The ouput of the critic network is a scalar.
@@ III-H Model-Predictive Value Expansion
In this section, we propose an extension to our algorithm that makes use of the predictions of the MPC to improve the quality of the value function estimate. We call this extension Model-Predictive Value Expansion (MPVE). This idea is inspired by the Model-Based Value Expansion algorithm proposed by [119], which uses the extra predicted rollouts using a learned model to improve the value function estimate. Assuming that our model is accurate up to depth $H$ and that we have access to the reward of the predictions, one can define the H-Step Model-Predictive Value Expansion as
@@
Figure 2: Actor-Critic Model Predictive Control (AC-MPC) applied to agile flight: velocity profiles and corresponding value function plots. The left side illustrates horizontal flight, while the right side shows vertical flight. In the value function plots, areas with high values (depicted in yellow) indicate regions with the highest expected returns. The MPC predictions are shown as black Xs.
@@ IV Experimental Setup
This section presents a set of experiments, both in simulation and the real world. All experiments have been conducted using a quadrotor platform. To showcase the capabilities of our method, we have chosen the task of agile flight through a series of gates in different configurations: horizontal, vertical, circular, and SplitS. Additionally, to show the sim-to-real transfer capabilities, both circular and SplitS tracks are deployed in the real world. We train in a simple simulator in order to speed up the training and evaluate in BEM, a high-fidelity simulator [120], which has a higher level of similarity in terms of aerodynamics with the real world. The quadrotor platform’s dynamics are the same as in [14]. For every different task, the policies are retrained from scratch. Lastly, all experiments have been conducted using a modification of the Flightmare software package [121] for the quadrotor environment and PPO implementation and Agilicious [122] for the simulation and deployment.
@@ IV-A1 Observation space
For all tasks presented in our manuscript, the observation space does not change, and it consists of two main parts: the vehicle observation $\mathbf{o}_{t}^{\text{quad}}$ and the race track observation $\mathbf{o}_{t}^{\text{track}}$. We define the vehicle observation as ${\mathbf{o}_{t}^{\text{quad}}=[\bm{v}_{t},\mathbf{R}_{t}]\in\mathbb{R}^{12}}$, which corresponds to the quadrotor’s linear velocity and rotation matrix. We define the track observation vector as $\mathbf{o}_{t}^{\text{track}}=[\delta\bm{p}_{1},\cdots,\delta\bm{p}_{i},\cdots],\;i\in[1,\cdots,N]$, where $\delta\bm{p}_{i}\in\mathbb{R}^{12}$ denotes the relative position between the vehicle center and the four corners of the next target gate $i$ or the relative difference in corner distance between two consecutive gates. Here $N\in\mathbb{Z}^{+}$ represents the total number of future gates. This formulation of the track observation allows us to incorporate an arbitrary number of future gates into the observation. We use $N=2$, meaning that we observe the four corners of the next two target gates. We normalize the observation by calculating the mean and standard deviation of the input observations at each training iteration.
@@ IV-A2 Action space
Previous work has demonstrated the high importance of the action space choice for learning and sim-to-real transfer of robot control policies [123, 124]. In this work, the control input modality is collective thrust and body rates, which was previously shown to perform best for agile flight [123]. This action is expressed as a 4-dimensional vector $\mathbf{a}=[c,\omega_{x},\omega_{y},\omega_{z}]\in\mathbb{R}^{4}$, representing mass-normalized thrust and body rates, in each axis separately. Even if the MPC block uses a model that limits the actuation at the single rotor thrust level, collective thrust and body rates are computed from these and applied to the system. This ensures that the computed inputs are feasible for the model of the platform.
@@ IV-A3 Rewards
For all experiments, one reward term in common is the gate progress reward, which encourages fast flight through the track. The objective is to directly maximize progress toward the center of the next gate. Once the current gate is passed, the target gate switches to the next one. At each simulation time step $k$, the reward function is defined by
@@ V Results
We perform multiple experiments to better understand different properties of our proposed approach. Namely, we study (i) AC-MPC’s performance on multiple drone racing tasks and its sample efficiency in comparison to AC-MLP policies, (ii) its robustness to external disturbances, (iii) its robustness to changes in the dynamics, (iv) the benefits of the Model-Predictive Value expansion, (v) the interpretability properties of AC-MPC, (vi) its sensitivity to exploration hyperparameters, (vii) its real-world deployment. Each aspect corresponds to a different research question that we address in each of the following sections.
@@ V-A Does our method improve performance and sample efficiency?
In Fig. 4, we can see the reward evolution for AC-MLP when compared with AC-MPC. This comparison was conducted after optimizing the initial standard deviation hyperparameter for both approaches and choosing the best result for each. Generally—as we also study in section V-F—AC-MPC performs better for lower exploration parameters, whereas AC-MLP needs higher exploration parameters. Fig. 4 shows that the asymptotic training performance and sample efficiency are improved when choosing AC-MPC, since it is able to leverage the prior information encoded in the dynamics within the differentiable MPC. This is particularly evident in the reward evolution observed for the Vertical and SplitS tracks. For simpler tasks, the prior knowledge introduced in AC-MPC is not strongly advantageous to performance. For instance, in the Horizontal track, which is relatively simple, the training performance of AC-MLP is on par with AC-MPC, since the maneuvers that the policy needs to discover to perform the task do not require an elaborate exploration behavior.
@@
Figure 4: Reward evolution for different tracks, for AC-MLP (for different N values of the horizon, $N=2$ and $N=5$) and AC-MPC. The values have been obtained after optimizing the initial exploration standard deviation for both approaches independently, and then selected the best result from each. For Vertical and SplitS tracks, one can observe how the AC-MPC approach is able to leverage its prior knowledge of the system and showcase improved learning. For the Horizontal track, because of its simplicity, the leverage of prior knowledge is diluted.
@@
Figure 5: Baseline comparisons between our AC-MPC, a standard PPO (termed AC-MLP), and a standard tracking MPC. (A): Robustness against wind disturbances (vertical track). All policies are trained without disturbances. Black arrows indicate the quadrotor’s attitude. (B): Robustness against changes in initial conditions (horizontal track). Trajectories are color-coded, with crashed trajectories in red and successful in blue.
@@
Figure 6: Success rate against variations of mass, inertia, and body rate limit. Policies are trained in the SplitS track with a nominal value for these parameters. This nominal value is depicted in red. Then, these policies are deployed in 64 parallel environments, each environment having slightly different initial position. In the deployment stage, mass, inertia, and body rate limit are varied. The success rate is computed as the proportion of trajectories that don’t have crashes. Because AC-MPC has the dynamics embedded in the network architecture, it shows to be more robust to these changes than its AC-MLP counterpart.
V-B Is our approach robust to external disturbances?
@@
We perform various studies where the standard actor-critic PPO architecture [14] (labeled as AC-MLP) and a standard tracking MPC are compared to our approach (labeled as AC-MPC) in terms of generalization and robustness to disturbances. AC-MLP and AC-MPC approaches are trained with the same conditions (reward, environment, observation, simulation, etc.). All these evaluations are conducted using the high-fidelity BEM simulator [120]. The MPC approach tracks a time-optimal trajectory obtained from [9].
@@
This is because, when using AC-MPC, we impose a modular dynamic structure compared to the flexibility of the single neural network used by the AC-MLP architecture.
In terms of disturbance rejection and out-of-distribution behavior, we conduct three ablations, shown in Fig. 5 and Table I). In Fig. 5A (and the Vertical Wind column of Table I), we simulate a strong wind gust that applies a constant external force of 11.5 N (equivalent to 1.5x the weight of the platform). This force is applied from $z=10m$ to $z=8m$. We can see how neither the AC-MLP nor the MPC policies can recover from the disturbance and complete the track successfully. On the other hand, AC-MPC achieves a higher success rate (83.33%, as shown in Table I), and exhibits more consistency among repetitions. This showcases that incorporating an MPC block enables the system to achieve robustness.
For the Vertical and Horizontal experiments in Table I, we simulate 10000 iterations for each controller where the starting points are uniformly sampled in a cube of 3m of side length where the nominal starting point is in the center. In the Horizontal case, the results are shown in Fig. 5B. It is important to highlight that during training of AC-MLP and AC-MPC, the initial position was only randomized in a cube of 1m of side length. The successful trajectories are shown in blue in Fig. 5B, while the crashed ones are shown in red. In Table I, we can observe that the AC-MPC presents a higher success rate than AC-MLP in both experiments. One can also see that AC-MPC has a higher success rate than the MPC approach in the Horizontal task, but this is not the case in the Vertical task. The reason behind this is that in the Vertical task, the MPC is not able to track the solution that turns the drone upside down, therefore resulting in the sub-optimal solution of setting all thrusts to near-zero state and dropping only by the effect of gravity, which results in slower but safer behavior. This is evident by looking at the average speed column.
@@
V-C Is our approach robust to changes in the dynamics?
Since our approach integrates a model-based controller that explicitly accounts for system dynamics, the question arises of whether AC-MPC is sensitive to dynamic parameters changes when compared to AC-MLP. To investigate this, we conduct a series of experiments in which key dynamic parameters are systematically varied. Both policies are trained on the SplitS track using nominal parameter values, after which they are evaluated across a range of perturbed parameters for both AC-MLP and AC-MPC. To quantify the robustness of these policies, we deploy them from 64 different positions, equally spread on a cube of $0.5\,m$ of side length. Success is defined as the percentage of trajectories that successfully navigate through all gates of the SplitS track. Fig. 6 presents the success rate as a function of variations in mass (up to $+27$%), in XY inertia (up to ±$\pm 80$%) , and body rate limits (up to ±$\pm 40\%$).

### v8

@@ III-B Tracking Model Predictive Control
The tracking MPC formulation also relies on the assumption that a trajectory represents a well-posed, quadratic proxy objective for solving the end task. While this assumption holds for many practical use cases, it becomes limiting in situations where the system operates near its performance limits, such as aggressive maneuvers or high-speed tasks, where the trajectory may no longer serve as an effective proxy. In such cases, tracking a optimal reference in the real world can result in suboptimal performance or even task failure [14].
@@ III-D General Quadratic MPC formulation
Instead of manually designing a complex, non-linear cost function that would subsequently require approximation, we propose to learn the final quadratic approximation directly. In this paper, we propose to directly search for the matrix coefficients of Eq. (4) (i.e., $Q_{k}$ and $p_{k}$) using reinforcement learning and differentiable MPC [31]. This approach allows the reinforcement learning agent to discover the most effective quadratic representation of the high-level task objective. By doing so, we can encode the task using a general, non-differentiable reward function while simultaneously producing a controller that is already computationally tractable, ready for real-time deployment without further approximations.
@@ III-E Actor-Critic Reinforcement Learning
where $\tau$ denotes a trajectory, $r(s_{k},a_{k})$ the reward function, and $\gamma\in[0,1)$ is a discount factor that discounts future rewards. Unlike real-time optimization methods like MPC, RL can handle non-differentiable, non-derivable models and easily optimize sparse rewards. Additionally, policy gradient optimization is typically performed offline through interactions with simulation. Once trained, the policy $\pi_{\theta}$ enables the computation of control signals by simply evaluating the function $a^{\ast}=\pi_{\theta}(s_{k})$ at each time step, reducing the computational complexity during deployment. For Actor-Critic reinforcement learning, the key idea is to simultaneously learn a state-value function $V_{\omega}(s)$ and learn a policy function $\pi_{\theta}$, where the value function (Critic) and policy (Actor) are parameterized by $\omega$ and $\theta$ separately. The policy is updated via policy gradient [121],
@@ III-F Actor-Critic Model Predictive Control
This paper proposes an Actor-Critic MPC controller architecture where the MPC is differentiable [31] and the cost function is learned end-to-end using RL. The MPC block is introduced as the differentiable module of the actor in an actor-critic pipeline that uses the proximal policy optimization (PPO) algorithm [122]. Our architecture is shown in Fig. 1. In contrast to previous work [64], where the MPC is taken as a black-box controller and the gradient is sampled, in our case the gradient of the cost function with respect to the solution is analytically computed and propagated using a differentiable MPC [31]. Therefore, for every backward and forward pass of the actor network, we need to solve an optimization problem. Instead of resorting to task-specific engineering of the cost function, we propose a neural cost map where the $Q_{k}$ and $p_{k}$ terms are the output of the neural network layers preceding the differentiable MPC block. This allows us to encode the end task directly as a reward function, which is then trainable end-to-end using the PPO training scheme. The main benefit of this approach with respect to training a standard Multi-Layer Perceptron (MLP) end-to-end is that the final module of the actor is a model-based MPC controller,
@@
and therefore it retains its generalizability to out-of-distribution situations. The model-based controller placed after the last layer of the actor network ensures that the commands are always feasible for the dynamics at hand, and that they respect the system constraints. This differentiable MPC supports control input constraints but not state constraints, which is why we add $\bm{u}_{k}\in\mathcal{U}$ in (1). To allow for exploration, during training the control inputs are sampled from a Gaussian distribution where the mean is the output of the MPC block, and the variance is controlled by the PPO algorithm. However, during deployment the output from the MPC is used directly on the system without further sampling, retaining all properties of a model-based controller. Algorithm 1 shows a high-level description of the AC-MPC algorithm.
@@ III-G Neural Cost Map
The purpose of the diagonalization of the $Q$ matrix is to reduce the dimensionality of the learnable parameter space. Therefore, the dimensionality of the output dimension of the Cost Map is $2T(n_{state}+n_{input})$. In order to ensure the positive semi-definiteness of the $Q$ matrix and the positive definiteness of the R matrix, a lower bound on the value of these coefficients needs to be set. To this end, the last layer of the neural cost map has been chosen to be a sigmoid which allows for upper and lower bounds on the output value. These lower and upper limits are chosen equal for $Q$ and $p$, of 0.1 and 100000.0, respectively. Therefore, the final neural cost map consists of two hidden layers of width 512 with ReLUs in between and a sigmoid non-linearity at the end. The critic network consists also of two hidden layers of width 512 and ReLUs. The output of the critic network is a scalar.
@@ III-H Model-Predictive Value Expansion
In this section, we propose an extension to our algorithm that makes use of the predictions of the MPC to improve the quality of the value function estimate. We call this extension Model-Predictive Value Expansion (MPVE). This idea is inspired by the Model-Based Value Expansion algorithm proposed by [123], which uses the extra predicted rollouts using a learned model to improve the value function estimate. Assuming that our model is accurate up to $H$ horizon steps and that we have access to the reward of the predictions, one can define the H-Step Model-Predictive Value Expansion as
@@
Figure 2: Actor-Critic Model Predictive Control (AC-MPC) applied to agile quadrotor flight: velocity profiles and corresponding value function plots. The left side illustrates horizontal flight, while the right side shows vertical flight. In the value function plots, areas with high values (depicted in yellow) indicate regions with the highest expected returns. The MPC predictions are shown as black Xs.
@@ IV Experimental Setup
This section presents a set of experiments, both in simulation and the real world. All experiments in this work are conducted on a quadrotor system, both in simulation and on a real platform. To showcase the capabilities of our method, we have chosen the task of agile flight through a series of gates in different configurations: horizontal, vertical, circular, and SplitS. Additionally, to show the sim-to-real transfer capabilities, both circular and SplitS tracks are deployed in the real world. We train in a simple simulator in order to speed up the training and evaluate in BEM, a high-fidelity simulator [124], which has a higher level of similarity in terms of aerodynamics with the real world. The quadrotor platform’s dynamics are the same as in [14]. For every different task, the policies are retrained from scratch. Lastly, all experiments have been conducted using a modification of the Flightmare software package [125] for the quadrotor environment and PPO implementation and Agilicious [126] for the simulation and deployment.
@@ IV-A1 Observation space
All tasks presented in our manuscript are quadrotor tasks, and the observation space has been tailored for such. The observation space for the quadrotor does not change, and consists of two main parts: the quadrotor’s state observation $\mathbf{o}_{t}^{\text{quad}}$ and the race track observation $\mathbf{o}_{t}^{\text{track}}$. We define the quadrotor’s state observation as $\mathbf{o}_{t}^{\text{quad}}=[\bm{v}_{t},\mathbf{R}_{t}]\in\mathbb{R}^{12}$}, which corresponds to the quadrotor’s linear velocity and rotation matrix. We define the track observation vector as $\mathbf{o}_{t}^{\text{track}}=[\delta\bm{p}_{1},\cdots,\delta\bm{p}_{i},\cdots],\;i\in[1,\cdots,G]$, where $\delta\bm{p}_{i}\in\mathbb{R}^{12}$ denotes the relative position between the vehicle center and the four corners of the next target gate $i$ or the relative difference in corner distance between two consecutive gates. Here $G\in\mathbb{Z}^{+}$ represents the total number of future gates. This formulation of the track observation allows us to incorporate an arbitrary number of future gates into the observation. We use $G=2$, meaning that we observe the four corners of the next two target gates. We normalize the observation by calculating the mean and standard deviation of the input observations at each training iteration.
@@ IV-A2 Action space
Previous work has demonstrated the high importance of the action space choice for learning and sim-to-real transfer of robot control policies [127, 128]. In our work, the quadrotor control input modality is collective thrust and body rates, which was previously shown to perform best for agile flight [127]. This action is expressed as a 4-dimensional vector $\mathbf{a}=[c,\omega_{x},\omega_{y},\omega_{z}]\in\mathbb{R}^{4}$, representing mass-normalized thrust and body rates, in each axis separately. Even if the MPC block uses a model that limits the actuation at the single rotor thrust level, collective thrust and body rates are computed from these and applied to the system. This ensures that the computed inputs are feasible for the model of the platform.
@@ IV-A3 Rewards
For all our experiments, the reward is tailored for quadrotor racing through gates. As such, one reward term in common is the gate progress reward, which encourages fast flight through the track. The objective is to directly maximize progress toward the center of the next gate. Once the current gate is passed, the target gate switches to the next one. At each simulation time step $k$, the reward function is defined by
@@ V Results
We perform multiple experiments to better understand different properties of our proposed approach. Namely, we study (i) AC-MPC’s performance on multiple drone racing tasks and its sample efficiency in comparison to AC-MLP policies, (ii) the choice of matrix parameterization, (iii) its behaviour against unknown external disturbances, (iv) its robustness to changes in dynamic parameters, (v) the benefits of the Model-Predictive Value expansion, (vi) the interpretability properties of AC-MPC, (vii) its sensitivity to exploration hyperparameters, (viii) its real-world deployment. Each aspect corresponds to a different research question that we address in each of the following sections.
@@ V-A Does our method improve performance and sample efficiency?
In Fig. 5, we can see the reward evolution for AC-MLP when compared with AC-MPC. This comparison was conducted after optimizing the initial standard deviation hyperparameter for both approaches and choosing the best result for each. Generally—as we also study in section V-G—AC-MPC performs better for lower exploration parameters, whereas AC-MLP needs higher exploration parameters. Fig. 5 shows that the asymptotic training performance and sample efficiency are improved when choosing AC-MPC, since it is able to leverage the prior information encoded in the dynamics within the differentiable MPC. This is particularly evident in the reward evolution observed for the Vertical and SplitS tracks. For simpler tasks, the prior knowledge introduced in AC-MPC is not strongly advantageous to performance. For instance, in the Horizontal track, which is relatively simple, the training performance of AC-MLP is on par with AC-MPC, since the maneuvers that the policy needs to discover to perform the task do not require an elaborate exploration behavior.
@@
Figure 4: Reward evolution for agile quadrotor flight in the Split-S track. Three different cost map representations are shown: Diagonal, Cholesky and Full Matrix.
@@
Figure 5: Reward evolution for drone racing in three different tracks, for AC-MLP (for different N values of the horizon, $N=2$ and $N=5$) and AC-MPC. The values have been obtained after optimizing the initial exploration standard deviation for both approaches independently, and then selected the best result from each. For Vertical and SplitS tracks, one can observe how the AC-MPC approach is able to leverage its prior knowledge of the system and showcase improved learning. For the Horizontal track, because of its simplicity, the leverage of prior knowledge is diluted.
@@
Figure 6: Baseline comparisons between our AC-MPC, a standard PPO (termed AC-MLP), a standard tracking MPC and a L1-MPC, for the task of agile quadrotor flight. (A): Robustness against wind disturbances (vertical track). All policies are trained without disturbances. Black arrows indicate the quadrotor’s attitude. (B): Robustness against changes in initial conditions (horizontal track). Trajectories are color-coded, with crashed trajectories in red and successful in blue.
@@
Figure 7: Success rate against variations of quadrotor’s mass, inertia, and body rate limit. Policies are trained in the SplitS track with a nominal value for these parameters. This nominal value is depicted in red. Then, these policies are deployed in 64 parallel environments, each environment having slightly different initial position. In the deployment stage, mass, inertia, and body rate limit are varied. The success rate is computed as the proportion of trajectories that don’t have crashes. Because AC-MPC has the dynamics embedded in the network architecture, it shows to be more robust to these changes than its AC-MLP counterpart.
V-B Does the choice of diagonal matrix affect performance?
In this section, we investigate a critical design choice within our framework: the structure of the learnable cost matrix. Our goal is to empirically demonstrate that constraining this matrix to be diagonal does not negatively impact performance. On the contrary, we aim to show that this simplification makes the learning problem more tractable.
To this end, we compare our standard diagonal representation of the quadratic cost matrix Q against two more complex alternatives: a full matrix decomposition and a Cholesky decomposition. All three methods produce the positive semi-definite matrices required for the optimization problem, but they differ in the number of learnable parameters, which directly influences the complexity of the learning task.
For the Cholesky case, the free parameters are the non-zero parameters of a triangular matrix $A$, where the diagonal elements are positive and the rest are free to be positive or negative. In this case, the number of free parameters would be $N(N+1)/2$. Similarly, for the full matrix case, the free parameters are all the elements of a matrix $A$, and therefore the number of free parameters would be $N^{2}$. In both cases, the matrix $Q$ is parameterized as follows:
$Q=AA^{T},$
This specific formulation is used to ensure that the resulting Q matrix is always positive semi-definite, a necessary condition for the stability of the underlying optimization problem.
Following this, we build three different cost map representations, mapping observations to cost function: Diagonal, Cholesky and Full Matrix. We train the three of them for three different tracks: Horizontal, Vertical and SplitS, and track the growth of the reward function to measure learning performance.
As shown in the learning curves in Fig. 4, the agent using the Diagonal representation is the only one capable of learning the required behavior, showing a clear and consistent increase in reward across all three tracks. In contrast, for both the Cholesky and Full Matrix cases, the network fails to learn within the same sample budget. We hypothesize that this outcome is a direct result of the higher learning difficulty imposed by the larger number of parameters in the more complex cost functions. This high-dimensional search space likely prevents the learning algorithm from finding a useful gradient, demonstrating that the constrained and more manageable problem offered by the diagonal matrix helps the training.
V-C Is our approach robust to external disturbances?
@@
We perform various studies where the standard actor-critic PPO architecture [14] (labeled as AC-MLP), a standard tracking MPC and a state-of-the-art adaptive L1-MPC [129] controller are compared to our approach (labeled as AC-MPC) in terms of generalization and robustness to disturbances. Both the AC-MLP and AC-MPC approaches are trained with the same conditions (reward, environment, observation, simulation, etc.). It is important to note that since we aim to test for out-of-distribution behaviour, the disturbances are not present at training time, only during testing. All these evaluations are conducted using the high-fidelity BEM simulator [124]. Both the MPC and the L1-MPC approaches track a time-optimal trajectory obtained from [9].
@@
In terms of disturbance rejection and out-of-distribution behavior, we conduct three ablations, shown in Fig. 6 and Table I). In Fig. 6A (and the Vertical Wind column of Table I), we simulate a strong wind gust that applies a constant external force of 11.5 N (equivalent to 1.5x the weight of the platform). This force is applied from $z=10m$ to $z=8m$. We can see how neither the AC-MLP nor the MPC policies can recover from the disturbance and complete the track successfully. The adaptive L1-MPC controller is able to overcome the disturbance and pass through the gates with a success rate of 41.66% and an average speed of 8.69 m/s.
On the other hand, AC-MPC achieves a higher success rate and speeds (83.33% and 10.76 m/s, as shown in Table I), and exhibits more consistency among repetitions. This showcases that incorporating an MPC block enables the system to achieve better out of distribution behaviour.
For the Vertical and Horizontal experiments in Table I, we simulate 10000 iterations for each controller where the starting points are uniformly sampled in a cube of 3m of side length where the nominal starting point is in the center. In the Horizontal case, the results are shown in Fig. 6B. It is important to highlight that during training of AC-MLP and AC-MPC, the initial position was only randomized in a cube of 1m of side length. The successful trajectories are shown in blue in Fig. 6B, while the crashed ones are shown in red. In Table I, we can observe that the AC-MPC presents a higher success rate than AC-MLP in both experiments. One can also see that AC-MPC has a higher success rate than both the MPC and the L1-MPC approaches in the Horizontal task, but this is not the case in the Vertical task. The reason behind this is that in the Vertical task, the MPC and L1-MPC are not able to track the solution that turns the drone upside down, therefore resulting in the sub-optimal solution of setting all thrusts to near-zero state and dropping only by the effect of gravity, which results in slower but safer behavior. This is evident by looking at the average speed column in Table I.
It is important to clarify that the tracking MPC approaches have difficulties in this task because the trajectories to be tracked are time-optimal trajectories. This means that they command the platform to go from hover to 100% thrust and body rates commands directly from the start. This makes these trajectories extremely difficult or impossible to track in the presence of any disturbance or model mismatch.
@@
V-D Is our approach robust to changes in the dynamics?
Our AC-MPC approach integrates a model-based controller that explicitly accounts for system dynamics, and therefore it benefits from the possibility of changing dynamic parameters without retraining, up to certain extent. The question arises of whether AC-MPC is less sensitive to dynamic parameters changes when compared to AC-MLP. To investigate this, we conduct a series of experiments in which key dynamic parameters are systematically varied. Both policies are trained on the SplitS track using nominal parameter values, after which they are evaluated across a range of perturbed parameters for both AC-MLP and AC-MPC. To quantify the robustness of these policies, we deploy them from 64 different positions, equally spread on a cube of 0.5m0.5\,m of side length. Success is defined as the percentage of trajectories that successfully navigate through all gates of the SplitS track. Fig. 7 presents the success rate as a function of variations in mass (up to +27+27%), in XY inertia (up to $\pm 80$%) , and body rate limits (up to $\pm 40\%$).